OpenStudy (anonymous):

Quick question: I am stumped for some reason my brain is not working today... How do I figure this out.. x+3/10=-3/5

OpenStudy (anonymous):

Assuming this is.. \[\frac{ (x+3) }{10 } = \frac{ -3 }{ 5 }\] Then cross multiply, to get... \[5(x+3) = -30\] Multiply the 5 through the bracket... \[5x + 15 = -30\] Subtract 15 off both sides... \[5x = -45\] Divide both sides by 5 \[x = -9\] As a check plug x = -9 into the original equation and you will see that indeed \[L.H.S = \frac{ (-9 + 3) }{ 10 } = \frac{ -6 }{ 10 } = \frac{ -3 }{ 5 } = R.H.S\] L.H.S is referring to the Left Hand Side of the equation, and R.H.S to Right Hand Side of the equation...

OpenStudy (kesumonu):

i think the person might be kidding with that question... anyway its good that someone has helped.

OpenStudy (anonymous):

Samuell explained it perfectly.